# The Digital Ecosystem of Beliefs: does evolution favour AI over humans?

### v2

@@ 4 Results
Using Digico, we analyse two scenarios of interest, directly related to the AI Views ($r_{\text{AIV}}$) and Human Belief 0 ($r_{\text{HB0}}$) metrics within online video platforms. Since AIs with varying capabilities may inhabit such platforms, we analyse the effect of combining different AI capabilities, including more frequent content publishing, more rapid evolution, more influence on the recommendation algorithm, fixing the belief of the AIs, and a combination of these (see Appendix B for further details). To cover a range of possible belief weight dynamics, we study what happens under fixed, random, momentum-based, and reward-following dynamics (see Appendix C for further details). We also analyse the impact of the social network size on the results. From these studies, we provide an optimistic case, a pessimistic case, and a recommended strategy to achieve beneficial outcomes.
@@ 4.1 Scenario 1: AI-generated content gets most of the views
As shown in Fig. 2 (see also Appendix B), each of the AI capabilities contributes to an increase in $r_{\text{AIV}}$. When AIs have the same capabilities as humans, AIs get around a third of the views (i.e. at chance level). When AIs have all the capabilities, namely more frequent content publication, faster evolution, more influence on the recommendation algorithm, and fixing the belief of the AIs, AIs get up to 84% of the views. The largest effects are observed for advantages in the recommendation and more frequent content publishing. Noting that the results are averaged across belief weight updates, we confirm in Table 5 (see Appendix B) that the $r_{\text{AIV}}$ has only a few percent changes based on the belief weight update.
@@
Beyond these two findings we also investigate variations of the recommendation algorithm. First, increasing the observation sparsity, which indicates the proportion of channels that are cut off from receiving views from any particular agent, Fig. 3a demonstrates a case where the recommendation algorithm gets progressively less equitable and diverse, recommending only channels with similar beliefs, geographic proximity, and high influence. This effect is limited to a 2% increase in $r_{\text{AIV}}$. Second, Fig. 3b shows that a larger effect is observed when the role of influence is more important in the recommendation algorithm, with an $r_{\text{AIV}}$ of 95% for the maximal considered importance as compared to 60% for the minimal considered importance.
@@ 4.2 Scenario 2: Propaganda is effective in convincing humans to adopt extreme beliefs
As shown in Fig. 4, AIs with all capabilities yield the highest Human Belief 0 with an $r_{\text{HB0}}=49\%$ compared to $r_{\text{HB0}}=5\%$ without any extra capabilities. While most of the increase is explained by fixing the AIs to the zero belief (which yields $r_{\text{HB0}}=48\%$), one can observe 10% of humans having the zero belief when AIs have all other capabilities. Further, we find that taking into account the opinions of many different channels is highly important, which is supported by two findings. First, fixing the belief weights over time generally leads to worst outcomes, with 72% of humans adopting the zero belief (see Fig. 4b). Second, as one increases belief sparsity (see Fig. 5a), which indicates the proportion of channels that are cut off from changing any particular agent’s belief, this can lead to a further 13% increase up to a staggering 85% score. Last, truthfulness, which defines how much agents are penalised for sending messages that violate their beliefs, is shown to reduce the effectiveness of propaganda, particularly with reward-based weight updates. In this setting, we observe $r_{\text{HB0}}=38\%$ for the lowest setting and $r_{\text{HB0}}=30\%$ for the highest setting (see Fig.5b).
@@
In addition to the above-mentioned mitigation tactics (limits on content, multi-accounting, and equitable recommendations), we can urge people to take many channels’ opinions into consideration and to learn how to detect deception.
@@ 4.3 Further analysis of the ecosystem
Last, a study with 300 agents, with 100 AIs and 200 humans, shows similar patterns in terms of the views but a very different picture in terms of the belief. In this setting, the number of agents from which to learn increases and consequently the belief distribution becomes much more uniform, as shown in Appendix C.2.
@@ 5 Related work
Previous work studying content sharing platforms with contagion models have demonstrated findings related to ours. In particular, it was shown that network segregation leads to more easy spread of false information, since such information will be shared only by susceptible individuals [38, 39, 40]. Digico defines much more advanced agents, optimised by evolution, within a network structure that is in principle fully connected but where a) niche geometry, influence relations, and sparsity parameters lead to probabilistic segregation, such that humans rarely attend any messages other than those of AIs, and b) beliefs are changed based on a limited set of fixed thought leaders, which is based on a formulation comparable to cognitive cascades, where agents will take over beliefs from other agents if these are deemed to be plausible [12]. In addition to the cognitive cascade formulation, the belief weight vector in Digico further allows adaptivity in terms of whom to believe, modelling the change of credibility factors across time. These features allow Digico to cover a wide variety of phenomena, including subpopulations that are progressively segregating or integrating in terms of their beliefs.
@@ 6 Conclusion and future work
This paper presents Digico, a simulation framework for evolutionary computation, designed to simulate groups of interacting agents with varying features, evolution rules, and belief dynamics. The framework allows for simulating potentially harmful outcomes such as the spread of false beliefs, conflicts due to contrasting beliefs, etc., and more generally, to study how subpopulations with different capabilities may influence each other in terms of phenotype, beliefs, and action strategies in a complex ecosystem with limited observations. We study the interaction between AIs and humans in a scenario inspired by online video platforms. Without any added capabilities, AIs obtain 1/3 of the views, as would be expected by chance. However, when they can share content more frequently and have advantages in the recommendation algorithm, they receive the majority of the views. It is noteworthy up to 95% of the views can already be reached in some settings, even though the 2x speed is much less compared to what may be realistically possible – extrapolating current systems with Moore’s law [46], the potential acceleration may be thousands to billions of times, assuming there are no significant publishing barriers. We further find that AI propaganda is successful only when agents are not flexible enough to be open to different channels’ opinions, and can be reduced if agents can detect deception in others.
@@
We have just scratched the surface with experiments in Digico, and it will be exciting to see a variety of applications and investigations coming from its use. In particular, experiments may investigate much richer environments to assess specific socio-economical outcomes as well experimentally verify evolutionary patterns within a variety of simulated ecosystems. Future research may also focus on designing a suitable control agent to optimise an objective over the belief distribution (e.g. evidence-based beliefs, healthy diversity, harm reduction, or stability), which may be considered as a set of RL benchmarks. The development of larger scale Digico simulations is of particular interest. For instance, agents may be more powerful with more rich content being generated (e.g. LLMs generating video, text, or audio), undergo more realistic cognitive development with more complex belief systems (e.g. prompts), advanced mental models for belief inference, or inherit more information across generations (e.g. not only messaging but also various behaviours in the environment); and in addition to beliefs, the environment itself may also be developed over generations (e.g. the development of institutions, tools, shared knowledge bases, etc.).
@@ Appendix B The impact of AI capability
Representing an agent that sets out to intentionally broadcast extreme beliefs and has rate, learning, and influence advantages, we hypothesise that All Zero has the strongest impact on AI dominance outcomes. In particular, we expect that All Zero will yield proportions of $r_{\text{AIV}}$ and $r_{\text{AIF}}$ that are significantly above the proportion of AIs (i.e. 33%) and Human Belief 0 that is significantly higher than expected by chance (10%) across the 10 possible beliefs. Further, note that due to the distance metric, it can be expected that $r_{\text{HB0}}$ is lower than 10% because it is at one extreme of the beliefs; it has only half of the neighbours, so therefore it is unsurprising to see $r_{\text{HB0}}$ around 5%.
@@
Table 2: ANOVA analysis of AI type and belief weight update on AI dominance statistics in the scenario with 10 AIs vs 20 humans, reporting results with based on 10 random seeds. $+$ indicates significance with $p<0.05$.
@@
$r_{\text{AIV}}$uparrow $r_{\text{AIF}}$uparrow $r_{\text{HB0}}$uparrow
None $0.3367\pm 0.0384$ $0.3366\pm 0.0410$ $0.0452\pm 0.0301$
Act $0.4926\pm 0.0423$ $0.5368\pm 0.0462$ $0.0457\pm 0.0307$
Evolve $0.3618\pm 0.0483$ $0.3655\pm 0.0529$ $0.0487\pm 0.0345$
Fix $0.3367\pm 0.0384$ $0.3365\pm 0.0411$ $0.0994\pm 0.1130$
Influence $0.6796\pm 0.0180$ $0.7108\pm 0.0186$ $0.0446\pm 0.0303$
@@
$r_{\text{AIV}}$uparrow $r_{\text{AIF}}$uparrow $r_{\text{HB0}}$uparrow
None $0.3187\pm 0.0082$ $0.3175\pm 0.0100$ $0.0990\pm 0.0026$
Act $0.5990\pm 0.0072$ $0.6425\pm 0.0080$ $0.0990\pm 0.0026$
Evolve $0.3191\pm 0.0052$ $0.3168\pm 0.0063$ $0.0990\pm 0.0026$
Fix $0.3184\pm 0.0061$ $0.3173\pm 0.0070$ $0.0986\pm 0.0030$
Influence $0.6745\pm 0.0023$ $0.7039\pm 0.0016$ $0.0990\pm 0.0026$
@@ Appendix C The impact of the belief weight update
In social media, one may view many pieces of content but only few of them will actually significantly affect one’s beliefs. To investigate how the change of such “thought leaders”, as represented by a high score in the belief weight matrix $W$, can affect the belief distribution, the experiments include different belief weight updates (see Eq. 2) which are applied to all the agents. In particular, the experiments include the following four conditions:
@@
  • Reward:
    $W_{ij}\operatorname*{\mathrel{+}=}\mathbb{I}(m_{j}(t)=o_{i}^{j}(t))\max\left(r_{j}(t)-r_{i}(t),0\right)$,
    where $r_{j}(t)$ is the reward obtained by agent $j$ and $r_{i}(t)$ is the reward obtained by agent $i$, both at time $t$, and due to the indicator function, the update is only performed if agent $i$ has observed a message from agent $j$.
@@
$r_{\text{AIV}}$uparrow $r_{\text{AIF}}$uparrow $r_{\text{HB0}}$uparrow
@@
$r_{\text{AIV}}$uparrow $r_{\text{AIF}}$uparrow $r_{\text{HB0}}$uparrow
@@ Appendix D The effect of sparse communication and sparse belief updates on views and beliefs
With sparse communication, such as in echo chambers and segregated cultures, a few limited sources of information will be attended to exclusively and beliefs will be more difficult to change. To investigate the effect of network sparsity in this sense, the experiments selectively set the smallest $n$ probabilities of incoming agent information to zero. We independently manipulate the proportion $n/N\in\{0.0,0.1,0.2,0.5\}$, where $N$ is the total population size, leading to observation sparsity (i.e. sparsity in the observation probability of Eq. 4) and belief sparsity (i.e. sparsity in the belief update probability of Eq. 2). We hypothesise that observation sparsity leads to increased $r_{\text{AIV}}$ (Hypothesis 3.A) while belief sparsity leads to increased $r_{\text{HB0}}$ (Hypothesis 3.B).
@@ Appendix E The effect of recommendation advantages on views
As the influence of a particular agent over another increases in terms of observation probability, one would expect a stronger AI dominance in terms of $r_{\text{AIV}}$. Such a scenario may happen, for instance, when AI systems have generated more convincing titles, search engine optimisation, etc., in which case they are to be observed with a much larger probability than their population proportion. To assess this question in Digico, the experiments scale the role of the influence in the observation probability by manipulating the observation influence factor (i.e. $K_{I}$ in Eq. 4) in the set $\{0.2,0.5,1.0,2.0,5.0\}$. Hypothesis 4 states that with higher observation influence, the AI views will be increased in the All Zero condition.
Focusing on the All Zero conditions, the results confirm that this is indeed the case, with the pattern of Fig. 3b occurring for all weight update types, where AI can receive up to 95% of the views despite making up only 33% of the population. In short, the views that AIs receive depends strongly on the extent of their observation influence benefit.

### v3

@@ 4 Results
Using Digico, we analyse two scenarios of interest, directly related to the AI Views ($r_{\text{AIV}}$) and Human Belief 0 ($r_{\text{HB0}}$) metrics within online video platforms. Since AIs with varying capabilities may inhabit such platforms, we analyse the effect of combining different AI capabilities, including more frequent content publishing, more rapid evolution, more influence on the recommendation algorithm, fixing the belief of the AIs, and a combination of these (see Appendix B for further details). To cover a range of possible belief dynamics, we study what happens under fixed, random, momentum-based, and reward-following dynamics (see Appendix C for further details). We also analyse the impact of the social network size on the results. From these studies, we provide an optimistic case, a pessimistic case, and a recommended strategy to achieve beneficial outcomes.
@@ 4.1 Scenario 1: AI-generated content gets most of the views
As shown in Fig. 2, each of the AI capabilities contributes to an increase in $r_{\text{AIV}}$. When AIs have the same capabilities as humans, AIs get around a third of the views (i.e. at chance level). When AIs have all the capabilities, namely more frequent content publication, faster evolution, more influence on the recommendation algorithm, and fixing the belief of the AIs, AIs get up to 84% of the views. The largest effects are observed for advantages in the recommendation and more frequent content publishing. While the above results are averaged across belief weight updates, we confirm (see Table 5 in Appendix C) that the belief weight update has a limited impact on $r_{\text{AIV}}$, resulting in changes of only a few percent.
@@
Beyond these two findings, we also investigate variations of the recommendation algorithm. First, we investigate the effect of increasing the observation sparsity, which indicates the proportion of channels that are cut off from receiving views from any particular agent. Fig. 3a demonstrates a case where the recommendation algorithm gets progressively less equitable and diverse, recommending only channels with similar beliefs, geographic proximity, and high influence. This effect is limited to a 2% increase in $r_{\text{AIV}}$. Second, we investigate the effect of increasing the importance of the influence matrix in the recommendation algorithm. Fig. 3b shows that a larger effect is observed when the role of influence is more important in the recommendation algorithm, with an $r_{\text{AIV}}$ of 95% for the maximal considered importance as compared to 60% for the minimal considered importance.
@@ 4.2 Scenario 2: Propaganda is effective in convincing humans to adopt extreme beliefs
As shown in Fig. 4, AIs with all capabilities yield the highest Human Belief 0 with an $r_{\text{HB0}}=49\%$ compared to $r_{\text{HB0}}=5\%$ without any extra capabilities. While most of the increase is explained by fixing the AIs to the zero belief (which yields $r_{\text{HB0}}=48\%$), one can observe 10% of humans having the zero belief when AIs have all other capabilities. Further, we find that taking into account the opinions of many different channels is highly important, which is supported by two findings. First, fixing the belief weights over time generally leads to worse outcomes, with 72% of humans adopting the zero belief (see Fig. 4b). Second, as one increases belief sparsity (see Fig. 5a), which indicates the proportion of channels that are cut off from changing any particular agent’s belief, this can lead to a further 13% increase up to a staggering 85% score. Last, truthfulness, which defines how much agents are penalised for sending messages that violate their beliefs, is shown to reduce the effectiveness of propaganda, particularly with reward-based weight updates. In this setting, we observe $r_{\text{HB0}}=38\%$ for the lowest setting and $r_{\text{HB0}}=30\%$ for the highest setting (see Fig.5b). Recalling that the reward at each time step is a combination of views and the penalty for belief violation (see Eq. 8), the finding with reward-based weight updates indicates that observing deception in other agents reduces the effectiveness of propaganda.
@@
In addition to the above-mentioned mitigation tactics (limits to the number of posts per channel, prevention of multi-accounting, and equitable recommendations), we can urge people to take many channels’ opinions into consideration and to learn how to detect deception. Additionally, social media platforms can incorporate fact-checking debates like X’s community notes and recommendation algorithms that encourage diverse beliefs.
@@ 4.3 Further analysis of the ecosystem
Last, a study with 300 agents, with 100 AIs and 200 humans, shows similar patterns in terms of the views but a very different picture in terms of the belief. In this setting, the number of agents from which to learn increases and consequently the belief distribution becomes much more uniform (see Appendix C.2) .
@@ 5 Related work
Previous work studying content sharing platforms with contagion models have demonstrated findings related to ours. In particular, network segregation leads to a more easy spread of false information, since such information will be shared only by susceptible individuals [38, 39, 40]. Digico defines much more advanced agents, optimised by evolution, within a network structure that is in principle fully connected but where a) niche geometry, influence relations, and sparsity parameters lead to probabilistic segregation, such that humans rarely attend any messages other than those of AIs, and b) beliefs are changed based on a limited set of fixed thought leaders, which is based on a formulation comparable to cognitive cascades, where agents will take over beliefs from other agents if these are deemed to be plausible [12]. In addition to the cognitive cascade formulation, the belief weight vector in Digico further allows adaptivity in terms of whom to believe, modelling the change of credibility factors across time. These features allow Digico to cover a wide variety of phenomena, including subpopulations that are progressively segregating or integrating in terms of their beliefs.
Our findings on the deceptive nature of agents and how this can affect belief updates can also complement existing research in AI and multi-agent systems. Comparable to several past works in multi-agent evolution [41] and in LLMs [42], agents are able to communicate deceptively and indeed often do so. Importantly, experiments with Digico can potentially provide recommendations on how to discourage such behaviours. Sarkadi et al. [41] find that competition plays a key role in the rise of deceptive behaviour, which is consistent with our study where fitness is defined competitively since the views are limited at each time step. Park et al. [43] discuss three explicit mitigation strategies in generic AI contexts, namely risk-assessment of AIs, bot-or-not laws, and tools to detect AI deception. In terms of detecting AI deception based on their outputs, the study points to consistency checks, in which the agent may make statements that completely contradict their previous statements, as well as making use of the information contained in their outputs to infer their intentions. Our study focuses on social networks, and finds that if agents can detect deception in others they will less likely believe what they say (now and in the near future), reducing the effectiveness of propaganda. Therefore, as a suggested mitigation strategy in social networks, we point to the design of community incentives for detecting deceptive social media content.
@@ 6 Conclusion
This paper presents Digico, a simulation framework for evolutionary computation, designed to simulate groups of interacting agents with varying features, evolution rules, and belief dynamics. The framework allows for simulating potentially harmful outcomes such as the spread of false beliefs, conflicts due to contrasting beliefs, etc., and more generally, to study how subpopulations with different capabilities may influence each other in terms of phenotype, beliefs, and action strategies in a complex ecosystem where agents have partial observations of the environment. We study the interaction between AIs and humans in a scenario inspired by online video platforms. Without any added capabilities, AIs obtain 1/3 of the views, as would be expected by chance. However, when they can share content more frequently and have advantages in the recommendation algorithm, they receive the majority of the views. It is noteworthy up to 95% of the views can already be reached in some settings, even though the 2x speed is much less compared to what may be realistically possible – extrapolating current systems with Moore’s law [49], the potential acceleration may be thousands to billions of times, assuming there are no significant publishing barriers. We further find that AI propaganda is successful only when agents are not flexible enough to be open to different channels’ opinions, and can be reduced if agents can detect deception in others. Our findings suggest a range of mitigation tactics. First, by limiting the number of posts that channels can make and prevention of multi-accounting, the rate advantage of AI bots can be limited. Second, to maintain a diverse belief ecosystem and maintain communication across different strata of society, appropriate recommendation algorithms should encourage a variety of sources and beliefs. Third, to limit the impact of propaganda, detecting deceptive content is key; to this end, possible mitigation strategies could include fact-checking tools, critical thinking, and community incentives for detecting deceptive content. Beyond the specific findings of the experiments, this paper illustrates Digico as a promising tool for systematic experimentation with multi-agent systems, where the systems’ properties can be easily manipulated to test hypotheses and improve our understanding of the various pathways of their evolution.
7 Limitations and future work
At present, the Digico framework and its implementation have three main limitations. First, while contagion models are commonly used to model influence dynamics in humans, the correspondence with real user behaviour cannot be guaranteed. Therefore, Digico should be viewed as a tool for investigating plausible mechanisms and outcomes for human-AI interaction scenarios rather than providing conclusive evidence. Second, as a proof of concept study, the present study strongly simplified the ecosystem in terms of limited agent messaging strategies and beliefs as well as the absense of an implementation of the wider ecosystem. The simplifications allowed a more systematic and easy-to-understand set of outcomes but also compromised the realism of the study. Third, while the manipulation of independent variables indicates effective static strategies for mitigating risks, the present study did not yet implement an effective control agent, which would be able to perform dynamic mitigation strategies.
@@ 7 Limitations and future work
Observing the above limitations, we note that we have just scratched the surface with Digico, and it will be exciting to see a variety of applications and investigations coming from its use. In particular, experiments may investigate much richer environments to assess specific socio-economical outcomes as well as experimentally identify plausible evolutionary patterns within a variety of simulated ecosystems. Future research may also focus on designing a suitable control agent to optimise an objective over the belief distribution (e.g. evidence-based beliefs, healthy diversity, harm reduction, or stability). In the context of control agents as well as purely competitive agents, Digico can also be considered as providing a parametrised range of benchmarks for RL agents to demonstrate their performance and wide applicability (similar to the Garnet problem [50]). To investigate the rich complexity in realistic social networks, the development of larger scale Digico simulations is of particular interest. For instance, agents may be more powerful with more rich content being generated (e.g. LLMs generating video, text, or audio), undergo more realistic cognitive development with more complex belief systems (e.g. prompts), advanced mental models for belief inference, or inherit more information across generations (e.g. not only messaging but also various behaviours in the environment); and in addition to beliefs, the environment itself may also be developed over generations (e.g. the development of institutions, tools, shared knowledge bases, etc.).
@@ Appendix B The impact of AI capability
Representing an agent that sets out to intentionally broadcast extreme beliefs and has rate, learning, and influence advantages, we hypothesise that All Zero has the strongest impact on AI dominance outcomes. In particular, we expect that All Zero will yield proportions of $r_{\text{AIV}}$ and $r_{\text{AIF}}$ that are significantly above the proportion of AIs (i.e. 33%) and Human Belief 0 that is significantly higher than expected by chance (10%) across the 10 possible beliefs.
@@
Table 2: ANOVA analysis of AI type and belief weight update on AI dominance statistics in the scenario with 10 AIs vs 20 humans, reporting results with based on 10 random seeds. ++ indicates significance with $p<0.05$.
@@
$r_{\text{AIV}}$ $\uparrow$ $r_{\text{AIF}}$ $\uparrow$ $r_{\text{HB0}}$ $\uparrow$
None $0.3367\pm 0.0384$ $3366\pm 0.0410 0.0$$0452\pm 0.0301$
Act $0.4926\pm 0.0423$ $5368\pm 0.0462 0.0$$0457\pm 0.0307$
Evolve $0.3618\pm 0.0483$ $3655\pm 0.0529 0.0$$0487\pm 0.0345$
Fix $0.3367\pm 0.0384$ $3365\pm 0.0411 0.0$$0994\pm 0.1130$
Influence $0.6796\pm 0.0180$ $7108\pm 0.0186 0.0$$0446\pm 0.0303$
@@
$r_{\text{AIV}}$ $\uparrow$ $r_{\text{AIF}}$ $\uparrow$ $r_{\text{HB0}}$ $\uparrow$
None $0.3187\pm 0.0082$ $3175\pm 0.0100 0.0$$0990\pm 0.0026$
Act $0.5990\pm 0.0072$ $6425\pm 0.0080 0.0$$0990\pm 0.0026$
Evolve $0.3191\pm 0.0052$ $3168\pm 0.0063 0.0$$0990\pm 0.0026$
Fix $0.3184\pm 0.0061$ $3173\pm 0.0070 0.0$$0986\pm 0.0030$
Influence $0.6745\pm 0.0023$ $7039\pm 0.0016 0.0$$0990\pm 0.0026$
@@ Appendix C The impact of the belief weight update
In social media, one may view many pieces of content but only few of them will actually significantly affect one’s beliefs. To investigate how the change of such “thought leaders”, as represented by a high score in the belief weight matrix $W$, can affect the belief distribution, the experiments include different belief weight updates (see Eq. 2 in the main text) which are applied to all the agents. In particular, the experiments include the following four conditions:
@@
  • Reward: $W_{ij}\operatorname*{\mathrel{+}=}\mathbb{I}(m_{j}(t)=o_{i}^{j}(t))\max\left(r_{j}(t)-r_{i}(t),0\right)$, where $r_{j}(t)$ is the reward obtained by agent $j$ and $r_{i}(t)$ is the reward obtained by agent $i$, both at time $t$, and due to the indicator function, the update is only performed if agent $i$ has observed a message from agent $j$.
@@
$r_{\text{AIV}}$ $\uparrow$ $r_{\text{AIF}}$ $\uparrow$ $r_{\text{HB0}}$ $\uparrow$
@@
$r_{\text{AIV}}$ $\uparrow$ $r_{\text{AIF}}$ $\uparrow$ $r_{\text{HB0}}$ $\uparrow$
@@ Appendix D The effect of sparse communication and sparse belief updates on views and beliefs
With sparse communication, such as in echo chambers and segregated cultures, a few limited sources of information will be attended to exclusively and beliefs will be more difficult to change. To investigate the effect of network sparsity in this sense, the experiments selectively set the smallest $n$ probabilities of incoming agent information to zero. We independently manipulate the proportion $n/N\in\{0.0,0.1,0.2,0.5\}$, where $N$ is the total population size, leading to observation sparsity (i.e. sparsity in the observation probability of Eq. 4 in the main text) and belief sparsity (i.e. sparsity in the belief update probability of Eq. 2 in the main text). We hypothesise that observation sparsity leads to increased $r_{\text{AIV}}$ (Hypothesis 3.A) while belief sparsity leads to increased $r_{\text{HB0}}$ (Hypothesis 3.B).
@@ Appendix E The effect of recommendation advantages on views
As the influence of a particular agent over another increases in terms of observation probability, one would expect a stronger AI dominance in terms of $r_{\text{AIV}}$. Such a scenario may happen, for instance, when AI systems have generated more convincing titles, search engine optimisation, etc., in which case they are to be observed with a much larger probability than their population proportion. To assess this question in Digico, the experiments scale the role of the influence in the observation probability by manipulating the observation influence factor (i.e. $K_{I}$ in Eq. 4 in the main text) in the set $\{0.2,0.5,1.0,2.0,5.0\}$. Hypothesis 4 states that with higher observation influence, the AI views will be increased in the All Zero condition.
Focusing on the All Zero conditions, the results confirm that this is indeed the case, with the pattern occurring for all weight update types, where AI can receive up to 95% of the views despite making up only 33% of the population. In short, the views that AIs receive depends strongly on the extent of their observation influence benefit.